A two Digit number is three less than seven times the sum of its digits if the digits are reversed, the new number is 18 less than the original number what is the original number?
Answer by mananth(16949)

let the number be xy
x in tens place
y in units place
..
10x+y=7(x+y)-3
10x+y=7x+7y-3
3x-6y=-3
x-2y=-1.............. equation 1
..
10y+x= 10x+y-18
9y-9x=-18
y-x=-2 ..............equation 2
..
Add equation 1 & 2
x-2y+y-x=-1-2
-y=-3
y=3
..
plug the value of y in equation 2
3-x=-2
-x=-5
x=5
the number is 53